OpenStudy (anonymous):

PLEASE HELP.. MEDAL ((&)) FAN !! Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 : |2x - 4| = 5 Equation #2 : |3x - 4| = -12 A) Equation #1 and Equation #2 have the same number of solutions. B) Equation #1 has more solutions than equation #2. C) Equation #1 has fewer solutions than equation #2. D) None of the statements above describe the number of solutions to equations shown

OpenStudy (anonymous):

2x-5=5 2x-5=-5 can you solve those 2 for me please

OpenStudy (anonymous):

Dont know how to but sorry im pressed for time w/ this question

OpenStudy (anonymous):

how do you think we would isolate x whats the first step

OpenStudy (anonymous):

by subtracting the 5 ?

OpenStudy (anonymous):

close to get the terms on the other side we do the reciprocal or opposite of the term so add the 5 then divide by 2

OpenStudy (anonymous):

x=? x=?

OpenStudy (anonymous):

-8?

OpenStudy (anonymous):

i got them messed up im sorry 2x-4=5 add the 4 divide by 2 2x-4=-5 add the 4 divide by 2

OpenStudy (anonymous):

3x-4=-12 add the 4 divide by 3 3x-4=12 add the 4 divide by 3 which one has more solutions

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

Is it the second one that has more solutions ,, 3x - 4 = 12 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So would it be C @KlOwNlOvE ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok thank you !